# TE-NeXt: A LiDAR-Based 3D Sparse Convolutional Network for Traversability Estimation

### v1

@@ Abstract
Giving an expert system the decision of where to move in a space makes the estimation of traversability a critical task, as there are a multitude of different scenarios. Therefore, it is not a trivial problem to develop models that guarantee an adequate generalisation at the extent of the particular problem to be solved. This paper presents TE-NeXt, a novel and efficient architecture for Traversability Estimation (TE) from sparse LiDAR point clouds based on a residual convolution block. TE-NeXt block fuses notions of current trends such as attention mechanisms and 3D sparse convolutions. TE-NeXt aims to demonstrate high capacity for generalisation in a variety of urban and natural environments, using well-known and accessible datasets such as SemanticKITTI, Rellis-3D and SemanticUSL. Thus, the designed architecture ouperforms state-of-the-art methods in the problem of semantic segmentation, demonstrating better results in unstructured environments and maintaining high reliability and robustness in urbans environments, which leads to better abstraction. Implementation is available in a open repository to the scientific community with the aim of ensuring the reproducibility of results.
@@ I Introduction
In a world where autonomous navigation is a reality, the correct estimation of traversability acquires the magnitude of an area of study of its own, as it takes a very important role in the control of the autonomous vehicle in many applications.
The theory of affordances, based on animal behavior and their surroundings, as outlined by [1], suggests that the environment provides perceptible cues or affordances that guide an observer’s actions. Similarly, extending this concept to robotics [2], the robot perceives the environment’s capability to be safely traversed, independently of the entity moving through it. These affordances might include features such as flat surfaces, clear paths, absence of obstacles, and stable ground.
While traversability estimation has now been widely recognised as a critical capability for mobile robots, it has historically focused on obstacle avoidance. In this traditional context, the primary goal of the robot is to avoid any physical contact with its surroundings and navigate only through open spaces, relying on measurements from proximity sensors to detect nearby objects [3]. Afterwards, numerous strategies were developed to ensure safe navigation by means of exteroceptive data; from the creation of alternative representations of space such as 2D [4, 5], 2.5D [6, 7] or occupancy grid maps that enable robot localisation, to the use of supervised learning algorithms, specifically, semantic segmentation on data extracted from RGB, RGB-D cameras [8, 9, 10, 11], as the traversability analysis problem can also be interpreted as a subclass of the generic multiclass semantic segmentation problem. Despite the limited robustness that these methods exhibit in various outdoor environmental contexts [12, 13], the former approach (extrapolated to a 3D domain) seems to be more suitable for traversability estimation. The evolution of sensor technologies in terms of affordability, accuracy and portability, is leading to a re-evaluation of the methodologies to assess traversability. Specifically, the combination of neural networks with LiDAR sensors has gained popularity. LiDAR’s immunity to changes in lighting conditions, unlike optical sensors as cameras, makes it particularly attractive for this purpose.
@@
Thus, in recent years on-road driving has received significant attention in the domain of semantic segmentation, providing databases that, although their primary purpose is not traversability estimation, can be easily adapted to the problem. Nevertheless, on-road driving does not reflect the full extent of the problem. Off-road environments pose distinct challenges due to their uneven terrain, pronounced textures, and irregular features. The presence of these unique hazards increases the need for the autonomous vehicle to understand its environment at a higher level. Thus, it raises the need to provide a method for several different environments.
This paper presents an environment-independent traversability estimation technique, specifically the design of a deep neural network based on the description of traversable and non-traversable areas regardless of the environment captured by a LiDAR sensor (Figure 1). Several improvements and modifications are proposed based on concepts stated from different aspects of deep learning, in particular convolutional networks and vision transformers (ViT). In addition, significant improvements in both macro and micro design have been carried out. In summary, the main contributions of this paper are:
  • TE-NeXt: A novel 3D sparse convolutional neural network architecture for Traversability Estimation. The architecture has been specifically designed and optimised for this problem.
@@
As a result, the proposed network is able to outperform other existing methods in unstructured environments, while maintaining robustness, high precision ratios.
The rest of the paper is organized as follows: Section II describes the current methodologies under which the problem of traversability estimation is addressed. Section III then details the proposed architecture for that purpose. Section IV focuses on the procedure performed in the different experiments and the results obtained. Finally, Section V draws the main conclusions.
@@ II RELATED WORK
This section provides an overview of the methods that constitute the literature on traversability estimation. It will mention the most relevant deep learning works under which the problem was originally addressed to the current methodologies employed.
The strong emergence of deep learning methods, particularly in the area of computer vision, brought about a paradigm shift, as contrary to traditional traversability estimation methods that based on hand-crafted features and heuristics, these new methods rely on raw sensor data, such as images or point clouds, to learn feature representations. In addition, they attracted a great deal of interest from the research community because of the more visual nature of the information obtained from the environment [16]. Therefore, appearance-based methods became one of the most widely used approaches to this task.
One of the first supervised methods was proposed in [17], where a Multi-Layer Perceptron (MLP) classifier combined with colour and texture features (extracted from RGB images) was used as a method for automatic classification of road and off-road terrain. However, the methods quickly shifted towards using visual information as input due to the results shown by convolutional networks in image classification tasks. In this context, network architectures such as GA-Nav [8] and DeconvNet [18] emerged. Both only use visual information as input for terrain classification to determine safe areas to traverse and both grouped the semantic segmentation into 6 and 5 classes respectively.
@@
Progress in high-resolution data processing [22] made it possible to extrapolate the concept of traversability, previously approached from a two-dimensional point of view, to a three-dimensional domain. Point clouds produced by LiDAR sensors have served as the basis for traversability estimation methods in a 3D environment [23, 24, 25], although pre-processing is required.
@@
A common practice in a three dimensional context is 3D voxel partition, in works such as [26, 27] the authors represent the point cloud as a 3D voxel occupancy map. This voxel representation avoids the need for commonly used elevation maps, which are prone to error in the presence of overhanging obstacles and in multi-storey or low ceiling scenarios. Apart from the internal representation of the environment, they proposed a 3D sparse encoder-decoder that exploits the sparsity of the voxel representation for efficient computation. However, there are different ways of representing point clouds, from [28] where they are arranged in a multilevel polar grid, which naturally takes into account the non-uniform density of point clouds, to cylindrical voxelisation [29] based on the cylindrical coordinate system, which adjusts to the variable dispersion of the LiDAR point cloud.
@@
As with 2D vision, methods that fused input information of different nature were developed in the 3D context. One of the most common proposals is the merging of surface normals extracted from LiDAR data with RGB images [30, 31, 32]. All of them utilize multi-modal fusion of RGB images and geometric information derived from LiDAR point clouds. The architectures proposed employ transformer-based encoder blocks to extract features from the input modalities. This allows them to leverage the advantages of transformers in modeling long-range dependencies and global context.
Similarly, other approaches have combined the point clouds with data collected from the interaction of the vehicle with the environment, not as a fusion of input data but to establish training labels. Works such as [33, 34, 35], automatically generate training labels from proprioceptive data. This kind of approach is known as self-supervised traversability, where labels are generated from past driving trajectories by labeling regions traversed by the vehicle as highly traversable. The wheel-contact points are projected into image coordinates and filtered for occlusions using LiDAR points. Unsupervised LiDAR denoising is performed for robust labeling in unstructured environments.
Finally, this paper presents TE-NeXt, an architecture based on a sparse variant of the U-Net neural network [36], modified and improved for its application to traversability and non-traversability verification by approaching it as a pointwise point cloud semantic segmentation problem. The proposed architecture requires a point cloud as input, as some of the methods mentioned in this section, and the feature extraction at different scales is given in the residual block designed taking notions from the two current trends such as vision Transformers and ConvNext [14] philosophy. TE-NeXt demonstrates robustness in structured environments, with performance in natural outdoor environments at the core of the network design, outperforming state-of-the-art methods in these environments.
@@ III TE-NeXt: AN ARCHITECTURE FOR TRAVERSABILITY ESTIMATION
TE-NeXt is a novel sparse convolutional network designed for 3D semantic segmentation of point clouds, with a specific focus on traversability estimation. TE-NeXt builds upon the U-Net style architecture [36], replacing the standard residual blocks, introduced by classical ResNet networks [37], with a customised block inspired by the ConvNeXt [14] design principles. What sets TE-NeXt apart from recent work in this domain is its aim to learn a more general, high-level understanding of the environment that is not tied to specific settings or robotic platforms.
@@ III-A Network Architecture Overview
As previously stated, the focus of recent works is on the identification of environmental attributes that facilitate the discernment of traversable from non-traversable areas. However, in other domains such as large-scale place recognition, which a priori generalisation is a fundamental requirement, the features are initiated as a vector of 1’s for those non-empty voxels [39, 40]. According to this view, the method presented in this paper continues such approach as it aims to optimise the abstraction process in every possible environment without favouring it in any of them.
Next, we provide a description of the TE-NeXt architecture, depicted in Figure 2. The proposed network adopts a hierarchical structure analogous to the U-Net [36], which is an encoder-decoder configuration, with the spatial dimensions initially reduced through strided convolution operations. This enlarges the receptive field size of the neurons at deeper layers, enabling them to capture a more extensive contextual information from the input. Following each of these operations, a normalisation layer, LayerNorm, and an activation function, ReLu, are employed for each level of network resolution. The latter is a common technique in convolutional networks to learn complex, non-linear mappings between inputs and outputs. Conversely, the utilisation of a normalisation technique that is not commonly employed in classical convolutional networks is justified by the promising outcomes observed in vision Transformers. The main reason why it has proved so effective is its invariance with respect to the batch size, since it is not usually possible to process large amounts of data simultaneously in a three-dimensional domain.
Subsequently, the spatial dimensions are incrementally augmented through upsampling operations, such as transposed convolutions. To compensate for the loss of spatial details during encoding, skip connections are employed to concatenate feature maps from the encoder path with those in the decoder path at the same resolution level. Unlike U-Net and applying the same policy of [26] reducing the number of skip connections to only the lower layers, it helps reduce false positive predictions, likely because the decoder does not have to predict many forwarded features from the encoder. Using fewer skip connections seems to act as a form of regularization and improves overall traversability prediction performance on its task. As a result, the neurons in the final layers of the decoder have a large effective receptive field, capturing both global context and local details. Finally, as a decoder, the network output holds the same dimensions as the input data, providing a probability value obtained via a Sigmoid function as the final layer.
@@ III-B TE-NeXt residual block
The core innovation in TE-NeXt lies in its custom residual block. This block draws inspiration from the design principles of ConvNeXt [14], a Convolutional Network architecture which has inspired numerous works that constitute the current state-of-the-art on free space detection in images [32]. It incorporates several key ideas from vision Transformers while maintaining the simplicity and efficiency of standard ConvNets. TE-NeXt aspires to learn rich, generalizable feature representations from large-scale 3D point cloud data. The residual block is despicted in Figure 3.
@@
Next, another linear convolution layer increases the channels back to the desired output. In essence, the final 1x1x1 convolution allows the residual block to: i) restore the original channel dimensionality after the bottleneck layers, ii) increase representational power, iii) perform projection mapping. Finally, if the input and output dimensions are the equal (N=M), the input is added to the output via a skip connection before the normalization layer, the regularization technique applied (drop Path [41]) and a Gaussian Error Linear Unit (GeLU)[42]. Regarding the activation function employed, GeLU is smoother than Rectified Linear Unit activation (ReLU) and is utilized in the most advanced Transformers. This skip connection allows gradients to flow more easily during backpropagation, mitigating the vanishing gradient problem in very deep networks. Furthermore, the residual block can learn an identity mapping when needed, making it easier to optimize.

### v2

@@ Abstract
Among the multitude of scenarios in which autonomous robots are intended to operate, natural environments present the most significant challenges in the context of traversability estimation. Conversely, although comparatively less complex, structured environments are the most prevalent in the broader sense of autonomous navigation due to the abundance of human-designed spaces, which offer more predictable features for navigation systems. Therefore, both of these considerations lead to the conclusion that there is a need to develop methods that are specifically tailored to unstructured environments, while ensuring that the performance in structured environments is not significantly compromised. Existing methods are vulnerable to the influence of stationary environmental conditions or are constrained by the platform or vehicle on which they are developed. In this context, this paper makes a contribution to the state of the art with TE-NeXt. This network is a customised and efficient architecture for traversability estimation from sparse LiDAR point clouds based on a encoder-decoder topology that includes several modifications regarding: (i) the input features; (ii) the structure of encoder-decoder resolution levels; and (iii) the constitution of the 3D convolutional block. The modifications provide the robot with a deeper understanding of traversability in natural and unstructured environments. Thus, the designed architecture presents a trade-off solution, exhibiting superior performance in unstructured environments (the Rellis-3D dataset) while maintaining high reliability and robustness in urban environments (the SemanticKITTI and the SemanticUSL dataset). The implementation of the presented approach is available at an open repository to ensure result reproducibility.
@@ I Introduction
In a world where autonomous navigation is a reality, the estimation of traversability has become a significant area of study in its own right. This is because it plays a crucial role in the control of autonomous vehicles in a wide range of applications.
While the estimation of traversability has now been widely acknowledged as a critical capability for mobile robots, it has historically been oriented towards the avoidance of obstacles. In this traditional context, the primary objective of the robot is to avoid any physical contact with its surroundings and to navigate exclusively through open spaces, relying on measurements from proximity sensors to detect nearby objects[1]. Subsequently, a multitude of strategies were devised to guarantee secure navigation through exteroceptive data. These included the development of alternative spatial representations, such as two-dimensional [2, 3] and 2.5D maps [4, 5].
Although this approach has demonstrated efficacy in structured contexts, the inherent limitations of these methodologies when deployed in diverse natural outdoor environments,[6, 7], in conjunction with the advancement of sensor technologies in terms of cost-effectiveness, precision, and portability are prompting a reassessment of the methodologies utilized to evaluate traversability in unstructured environments. The natural environments present a series of distinctive challenges, largely due to the presence of uneven terrain, pronounced textures, and irregular features. In light of these unique hazards, there is a growing necessity for autonomous vehicles to achieve a more comprehensive understanding of their surrounding environment.
@@
The foundation for traversability estimation in unstructured environments was laid by pioneering programs such as DEMO III [8], PerceptOR [9] and DARPA grand challenge, which encouraged the development of innovative approaches to perception, learning, and control systems for autonomous ground vehicles navigating challenging off-road terrains. The resulting methodologies established three primary approaches: pure geometric perspective, geometric and vision-based methods, and vehicle-terrain interaction-based techniques. However, the landscape of traversability estimation has evolved significantly with the emergence of deep vision technologies. This advancement has led to a decline in purely geometric approaches and highlighted limitations in methods based on vehicle-terrain interactions, particularly due to the difficulty in obtaining consistent negative examples. As a result, the field has shifted towards more sophisticated, geometric and 3D vision-based techniques that can better address the complexities of traversability estimation in natural environments.
To enhance the efficacy of specialised techniques for off-road traversability estimation, this study proposes TE-NeXt, a LiDAR-based 3D sparse convolutional architecture capable of discerning traversable and non-traversable areas captured by a LiDAR sensor (Figure 1), enriching the perception of the environment with semantic information in addition to geometric information. In summary, the main contributions of this paper are:
  • TE-NeXt: A customised 3D sparse convolutional neural network architecture for Traversability Estimation. The architecture has been specifically designed and optimised for this problem.
@@
  • Applicability of the proposed method in an autonomous navigation system in dynamic environments.
Consequently, the proposed network is capable of outperforming existing methods in unstructured environments while maintaining robustness and high precision ratios in urban environments.
The remaining sections of the paper are structured as follows: Section II outlines the prevailing methodologies employed in addressing the issue of traversability estimation. Section III presents the proposed architecture in detail. Following, Section IV describes an exhaustive set of experiments carried out on different datasets with the aim of comparing the current approach with other cutting-edge methods.
@@ II RELATED WORK
This section provides an overview of the methods that constitute the literature on traversability estimation. It will mention the most relevant deep learning works under which the problem was originally addressed, as well as the current methodologies employed.
II-A Vision-based traversability estimation
The advent of deep learning methods, particularly in the domain of computer vision, has precipitated a paradigm shift. In contrast to traditional traversability estimation methods, which are based on hand-crafted features and heuristics, these new methods rely on raw sensor data, such as images or point clouds, to learn feature representations. Furthermore, they attracted considerable interest from the research community due to the more visual nature of the information obtained from the environment [12]. Consequently, appearance-based methods became one of the most widely used approaches to this task.
One of the earliest supervised methods was proposed in [13], where a Multi-Layer Perceptron (MLP) classifier combined with colour and texture features (extracted from RGB images) was used as a method for automatic classification of road and off-road terrain. However, the methods rapidly evolved towards the utilisation of visual information as an input, largely due to the impressive results demonstrated by convolutional networks in image classification tasks. In this context, network architectures such as GA-Nav [14] and DeconvNet [15] emerged. Both solely employ visual information as an input for terrain classification, with the objective of identifying safe traversal routes. Additionally, both approaches employed semantic segmentation, grouping the results into six and five classes, respectively.
@@ II-B Geometric and hybrid methods for traversability estimation
The advancement of high-resolution data processing techniques [19] has enabled the extrapolation of the concept of traversability, which was previously approached from a two-dimensional perspective, to a three-dimensional domain. Point clouds produced by LiDAR sensors have served as the basis for traversability estimation methods in a three-dimensional environment [20, 21, 22], although pre-processing is required.
@@
A common practice in a three dimensional context is 3D voxel partition, in approaches such as [23, 24] the authors represent the point cloud as a 3D voxel occupancy map. This voxel representation avoids the need for commonly used elevation maps, which are susceptible to inaccuracies in the presence of overhanging obstacles and in multi-storey or low ceiling scenarios. In addition to the internal representation of the environment, they proposed a 3D sparse encoder-decoder that exploits the sparsity of the voxel representation for efficient computation. However, there are various methods of representing point clouds. An approach that is outlined in [25], involves the arrangement of point clouds within a multilevel polar grid, which naturally takes into account the non-uniform density of point clouds. An alternative approach is presented in [26], where point clouds are represented using cylindrical voxels based on the cylindrical coordinate system.
@@
As with 2D vision, methods that fused input information of different nature were developed in the 3D context. One of the most common proposals is the merging of surface normals extracted from LiDAR data with RGB images [27, 28, 29]. All these studies utilise multi-modal fusion of RGB images and geometric information derived from LiDAR point clouds. The proposed architectures employ transformer-based encoder blocks to extract features from the input modalities.
TABLE I: Overview of related works.
Reference Input data Method Description Platform-specific
Selvathai et al., 2017 [13] RGB images MLP
Terrain classification into road and off-road
categories.
No
Guan et al., 2022 [14] RGB images
CNN-Attention Module
Deep Reinforcement Learning
Segmentation of different types of terrains
based on their navigability.
Yes
Sharma et al., 2017 [15] RGB images CNN
A semantic segmentation method for off-road
autonomous driving using transfer learning.
No
Valada et al., 2016 [16]
RGB,
Infrared images
CNN
Early and late fusion comparison with
images of differing natures.
No
Liu et al., 2023 [17] Infrared images CNN
A semantic segmentation method for
infrared road detection.
No
Xiong et al., 2021 [18] Thermal images
CNN-Attention Module
Thermal image semantic segmentation of nighttime driving scenes. No
Bae et al., 2023 [20] Point clouds
CNN-Clustering
Binary segmentation method for traversable and non-traversable regions. No
Lee et al., 2021 [21] Elevation maps MLP Method employs the extraction of terrain features to generate a traversability map. Yes
Seo et al., 2023 [22] Point clouds
CNN-MLP
The method employs proprioceptive sensor to infer binary and continuous traversability. No
Frey et al., 2022 [23] 3D maps Sparse CNN The method compute 3D occupancy traversa- bility map from data generated in simulation. No
Ruetz et al., 2024 [24] 3D maps Sparse CNN The method compute 3D occupancy traversa- bility map from terrain features. No
Fusaro et al., 2023 [25] Point clouds SVM The method operates on a pyramid-polar space representation to infer binary traversability. No
Min et al., 2022 [27]
RGB images,
Surface Normal
CNN-Attention Module
MLP
Off-road freespace detection method on the RGB image plane. No
Ye et al., 2023 [28]
RGB images,
Surface Normal
CNN-Attention Module
MLP
A multi-modal network for freespace detection. No
Li et al., 2024 [29]
RGB images,
Surface Normal
CNN-Attention Module
MLP
A novel Transformer architecture for semantic road scene parsing. No
Elnoor et al., 2024 [30]
Proprioceptive signals
PCA A method that uses proprioceptive data to evaluate terrain’s traversability in real time. Yes
Seo et al., 2023 [31]
Point clouds,
RGB images
CNN Method that learns traversability only from self-supervisions without explicit labels. No
Wellhausen et al., 2019 [32]
RGB images,
Torque Signal
CNN The method derives a ground reaction score, which serves as a measure of traversability. . Yes
Seo et al., 2023 [33]
Point clouds,
IMU
Sparse CNN The method predicts traversability cost map derived from vehicle-terrain interactions. Yes
Gasparino et al., 2022 [34]
RGB images,
Depth image
CNN Traversability prediction in image space pro- vides traversability coefficient for trajectories. Yes
Schreiber et al., 2024 [35] RGB images Mean Teacher-CNN Method of relative traversability estimation that uses sparsely annotated point pairs. No
Gao et al., 2021 [36] RGB images
CNN,
Clustering
Traversability is predicted by comparing the features to the clustered features. No
Ours Point clouds
Sparse CNN
A point-wise semantic segmentation method for traversability estimation. No
II-C Self-supervised traversability estimation
Similarly, as the primary drawback of the supervised methodology is the time-consuming annotation procedure, alternative techniques have been developed with the objective of reducing the labelling process. Consequently, methodologies such as that described in [30] have been proposed. These methods utilise statistical techniques, including principal component analysis (PCA), to estimate the traversability of a given surface based on the analysis of proprioceptive sensor signals. Furthermore, methodologies such as those described in [33, 32, 34, 31] automatically generate training labels from proprioceptive data in conjunction with image or LiDAR data, where labels are generated exploiting a vehicle’s driving experience from past trajectories.
However, the inability of these methods to achieve negative examples, i.e. moving through totally untraversable areas, without compromising the structure of the vehicle or robot, has prompted the development of alternative annotation processes. In solutions such as [35], relative annotations for pairs of points are supplied, indicating which point in the pair is more traversable, or if both points possess equal traversability. Alternative approaches, such as referenced in [36], employ sparse anchor patches to denote regions with distinct semantic attributes regarding their traversability and classify them into existing semantic clusters.
II-D Current proposal
This paper presents TE-NeXt, an encoder-decoder architecture optimised for traversability estimation in unstructured environments. The approach presented here can be classified as a supervised method, in accordance with the paradigm described in Subsection II-B, as it is a supervised paradigm method. However, it employs local point clouds, a distinctive feature that sets it apart from other methods, such as those described in [23, 24], which utilise point cloud maps and preprocess the data. Moreover, in contrast to the techniques outlined in Subsection II-A, TE-NeXt employs an optical sensor that is invariant to changes in light conditions (LiDAR sensor). Furthermore, the proposed methodology is designed to overcome the limitations associated with self-supervised learning approaches. Unlike existing methods, it does not require a specific platform, making it a more general and secure solution for all types of vehicles. Table I outlines the distinctive features of the proposed approach in comparison to existing, state-of-the-art methodologies.
The proposed architecture requires a point cloud as input, as some of the methods mentioned in this section, and the feature extraction at different scales is computed at the residual block, which inherits design notions from two broad paradigms: vision Transformers and ConvNext [10]. The primary objective of TE-NeXt is to demonstrate enhanced performance in natural environments in comparison to existing state-of-the-art methodologies for semantic segmentation and traversability estimation. Additionally, the system is designed to operate as a viable alternative in structured environments, although this is not the ultimate intended outcome of the study.
@@ III TE-NeXt: AN ARCHITECTURE FOR TRAVERSABILITY ESTIMATION
TE-NeXt is a dedicated convolutional neural network, specifically designed for the task of 3D semantic segmentation of point clouds, with a particular emphasis on the estimation of traversability. TE-NeXt builds upon the U-Net style architecture [37], replacing the standard residual blocks, introduced by classical ResNet networks [38], with a propietary block inspired by the ConvNeXt [10] design principles. What sets TE-NeXt apart from recent work in this domain is its aim to learn a more general, high-level understanding of the environment that is not tied to specific settings or robotic platforms.
@@ III-A Network Architecture Overview
As previously stated, the focus of recent works is on the identification of environmental attributes that facilitate the discernment of traversable from non-traversable areas. However, in other domains such as large-scale place recognition, which a priori generalisation is a fundamental requirement, the features are initiated as a vector of 1’s for those non-empty voxels [40, 41]. According to this view, the method presented in this paper continues such approach as it aims to optimise the abstraction process.
Next, we provide a description of the TE-NeXt architecture, depicted in Figure 2. The proposed network adopts a hierarchical structure analogous to the U-Net [37], which is an encoder-decoder configuration, with the spatial dimensions initially reduced through strided convolution operations. This enlarges the receptive field size of the neurons at deeper layers, enabling them to capture a more extensive contextual information from the input. Following each of these operations, a normalisation layer, LayerNorm, and an activation function, ReLu, are employed for each level of network resolution. The latter is a common technique used in convolutional networks to learn complex, non-linear mappings between inputs and outputs. Conversely, the utilisation of a normalisation technique that is not commonly employed in classical convolutional networks is justified by the promising outcomes observed in vision Transformers. The principal reason for its efficacy is its invariance with respect to the batch size, given that it is not typically feasible to process substantial quantities of data in a three-dimensional domain concurrently.
Subsequently, the spatial dimensions are incrementally augmented through upsampling operations, such as transposed convolutions. To compensate for the loss of spatial details during encoding, skip connections are employed to concatenate feature maps from the encoder path with those in the decoder path at the same resolution level. Unlike U-Net and applying the same policy of [23] reducing the number of skip connections to only the lower layers, it helps reduce false positive predictions, likely because the decoder does not have to predict many forwarded features from the encoder. A reduction in the number of skip connections appears to work as a form of regularisation, thus enhancing overall traversability prediction performance.
As a result, the neurons in the final layers of the decoder have a large effective receptive field, capturing both global context and local details. Finally, as a decoder, the network output holds the same dimensions as the input data, providing a probability value obtained via a Sigmoid function as the final layer.
@@ III-B TE-NeXt residual block
A major innovation of TE-NeXt can be attributed to its residual block. This block is inspired by the design principles of ConvNeXt [10], a convolutional network architecture that has informed numerous methods which constitute the current state-of-the-art on free space detection in images [29]. It incorporates several key ideas from Vision Transformers while maintaining the simplicity and efficiency of standard Convolutional Neural Networks (CNNs). TE-NeXt aims to learn rich, generalisable feature representations from large-scale 3D point cloud data. The residual block is illustrated in Figure 3.
@@
Next, another linear convolution layer increases the channels back to the desired output. In essence, the final 1x1x1 convolution allows the residual block to: i) restore the original channel dimensionality after the bottleneck layers; ii) increase representational power and iii) perform projection mapping. Finally, if the input and output dimensions are the equal (N=M), the input is added to the output via a skip connection before the normalization layer, the regularization technique applied (drop Path [42]) and a Gaussian Error Linear Unit (GeLU)[43]. Regarding the activation function employed, GeLU is smoother than Rectified Linear Unit activation (ReLU) and is utilized in the most advanced Transformers. This skip connection allows gradients to flow more easily during backpropagation, mitigating the vanishing gradient problem in very deep networks. Furthermore, the residual block can learn an identity mapping when needed, making it easier to optimize.